### v1

# Balancing Computation Load and Representation Expressivity in Parallel Hybrid Neural Networks

Abstract
Attention and State-Space Models (SSMs) when combined in a hybrid network in sequence or in parallel provide complementary strengths. In a hybrid sequential pipeline they alternate between applying a transformer to the input and then feeding its output into a SSM. This results in idle periods in the individual components increasing end-to-end latency and lowering throughput caps. In the parallel hybrid architecture, the transformer operates independently in parallel with the SSM, and these pairs are cascaded, with output from one pair forming the input to the next. Two issues are (i) creating an expressive knowledge representation with the inherently divergent outputs from these separate branches, and (ii) load balancing the computation between these parallel branches, while maintaining representation fidelity. In this work we present FlowHN, a novel parallel hybrid network architecture that accommodates various strategies for load balancing, achieved through appropriate distribution of input tokens between the two branches. Two innovative differentiating factors in FlowHN include a FLOP aware dynamic token split between the attention and SSM branches yielding efficient balance in compute load, and secondly, a method to fuse the highly divergent outputs from individual branches for enhancing representation expressivity. Together they enable much better token processing speeds, avoid bottlenecks, and at the same time yield significantly improved accuracy as compared to other competing works. We conduct comprehensive experiments on autoregressive language modeling for models with 135M, 350M, and 1B parameters. FlowHN outperforms sequential hybrid models and its parallel counterpart, achieving up to 4× higher Tokens per Second (TPS) and 2× better Model FLOPs Utilization (MFU).
1 Introduction
Transformer-based language models have become the foundation of modern Natural Language Processing (NLP) thanks to their ability to learn rich contextual representations through their self-attention mechanism for capturing fine-grained token interactions. While long-range relationships can also be captured, the quadratic cost of self-attention, requiring pairwise token interactions across the entire context grows prohibitively expensive in both computation and memory as sequence lengths increase (Zeng et al. (2025)). To address the limitations of traditional attention mechanisms in handling long sequences, researchers have proposed various alternatives, including Linear Attention (Katharopoulos et al. (2020)), Gated Linear Attention (Yang et al. (2023)), and more recently, State Space Models (SSMs) (Fu et al. (2022); Gu and Dao (2023)). SSMs offer constant-time recurrence and hardware-friendly implementations, making them efficient for long-range dependencies. However, while SSMs excel in modeling long range interactions with low computational overhead, they often under-perform compared to Transformers on tasks requiring fine-grained token interactions.
Figure 1: The overall architecture of our proposed parallel hybrid model, FlowHN
Hybrid networks have thus been proposed to combine the strengths of both, attention and SSM mechanisms to capture both short and long range dependencies. The two can be operative in sequence or in parallel. Our work presented in this paper is in the parallel category. Hymba (Dong et al. (2024)) is the only other parallel hybrid model that we are aware of with the same objective. The novelty in our proposed architecture lies in two important aspects. The first is a method for dynamic load balancing between the attention and SSM sub-blocks through appropriate distribution of input tokens with low information loss. The second novelty is a method to integrate the inherently divergent information coming independently from the two branches, reduce the information loss caused by independent parallel processing and yet capture knowledge effectively. These methods will be described in detail in later sections. We call this hybrid network as FlowHN.
Our main contributions are:
  • A novel parallel block (Figure 1) that processes each input through two concurrent branches, an attention branch for capturing global context and a SSM branch for memory-efficient, long-range dependency capture. The inherently divergent outputs from these concurrent branches are fused and projected to obtain a more expressive combined representation that forms the input to the next parallel block.
  • We present three token distribution strategies, including a dynamic FLOP-aware circulating split algorithm that adaptively distributes tokens between the two branches within each parallel block, thus balancing workload between branches and reducing total compute time with only slight information loss as compared to no splitting.
  • Through extensive experiments on standard text benchmarks and across three model scales (135M, 350M, and 1B parameters), we demonstrate that our architecture consistently surpasses both its parallel counterpart and leading sequential architectures in accuracy and computational efficiency.
2 Related Works
Large-scale autoregressive language models have demonstrated remarkable sequence processing capabilities, ranging from zero and few shot learning to long-range reasoning by pretraining on massive text corpora. However, their reliance on quadratic self-attention, which scales as $O(N^{2})$ in sequence length, imposes steep compute and memory demands that hinder processing of long contexts (Katharopoulos et al. (2020)). This challenge has spurred two complementary lines of research to reduce/eliminate self-attention compute costs, subquadratic attention approximations and state-space models, each with their individual strengths.
Subquadratic Sequence Modeling. Early work on reducing the quadratic bottleneck reformulated self-attention as a linear kernelized dot-product, enabling $O(N)$ complexity via associativity of matrix products (e.g., Linear Attentions, Katharopoulos et al. (2020)). Concurrently, RNN-inspired architectures such as RWKV blend parallelizable training with recurrent inference to achieve subquadratic scaling without softmax attention (Peng et al. (2023)). These approaches maintain competitive performance on long-sequence tasks while dramatically cutting FLOPs and memory usage. Yet they may struggle with some problems which full attention transformers are good at, such as content based reasoning. Although subquadratic sequence models dramatically reduce FLOPs and memory, they still struggle to preserve high-fidelity and long-range dependencies (Han et al. (2024b)) and suffer from representation collapse due to the non-injective nature of linearized attention kernels, which can assign identical weights to different queries (Han et al. (2024a)).
State-Space Models. Recently attention-free sequence modeling has centered on SSM architectures, which efficiently capture long-range dependencies without the quadratic cost of self-attention. The Structured State Space Sequence (S4) model (Gu et al. (2021)) introduced a novel parameterization of the linear SSM kernel that yields fast convolutional approximations and achieves competitive performance on long sequences. Building on S4, the H3 architecture (Fu et al. (2022)) interleaves shift and diagonal SSM matrices with multiplicative input projections, substantially narrowing the expressivity gap between pure SSMs and Transformer layers. Mamba (Gu and Dao (2023)) further unifies these selective SSM blocks with standard MLP layers, employing a selective-scan algorithm and hardware-aware optimizations (parallel scan, kernel fusion, and recomputation) to achieve true linear-time scaling and up to 5x higher TPS (Tokens/sec) than comparable Transformer-based models across language, audio, and genomics benchmarks. More recently, Mamba-2 (Dao and Gu (2024)) leverages structured state-space duality (SSD) to simplify parameter generation, producing SSM coefficients at block inception rather than as input-dependent functions, yielding 2–8x speedups over the original Mamba while delivering superior modeling quality relative to Transformer baselines.
Sequential Hybrid Models. Studies have highlighted that purely linear SSMs exhibit limited retrieval capabilities and thus benefit from integrating attention mechanisms to restore precise memory access (Jelassi et al. (2024); Abdin et al. (2024); Mohtashami and Jaggi (2023)). MambaFormer (Park et al. (2024)) extends this hybridization by replacing MLP blocks in standard Transformers with selective Mamba SSM blocks, thereby leveraging recurrent state updates to enhance long-range in-context learning and retrieval tasks. Jamba (Lieber et al. (2024); Team et al. (2024)) adopts a mixture-of-experts (MoE) design, interleaving Transformer, Mamba, and MoE layers to achieve high TPS, reduced memory footprint, and robust performance on contexts up to 256K tokens. SAMBA (Ren et al. (2024)) presents a lightweight, layer-wise fusion of selective SSM (Mamba) with the Sliding Window Attention technique, compressing distant context into recurrent hidden states while preserving fine-grained local attention, thus enabling efficient modeling of effectively unlimited sequences. Despite these advances, existing sequential hybrids are unable to avoid the bottlenecks of sequential stacking and thus do not efficiently exploit both attention and recurrence.
Parallel Hybrid Models. In Dong et al. (2024) Hymba, a family of small language models featuring a hybrid-head parallel architecture is introduced. To further bolster contextual recall, Hymba makes use of learnable meta tokens prepended to input sequences, which serve as a shared, dynamic cache across attention and SSM branches, thereby concentrating model capacity on salient information. The hybrid-head module is also optimized via a mix of global and partial sliding-window local attention, combined with cross-layer key–value cache sharing to sharply reduce memory overhead without compromising recall fidelity. However, despite these enhancements, Hymba faces computational efficiency challenges: the two branches process inputs at different speeds, causing the faster branch to idle while waiting for the slower one, ultimately leading to TPS bottlenecks. Another disadvantage is that the use of learnable meta tokens when concatenated to the input sequence increases the sequence length, which permeates through all stages of the pipeline and thus introduces additional computation/memory overheads.
In comparison, our FlowHN architecture balances the computational work load across attention and SSMs more evenly, does not require learnable meta tokens, so processing of extraneous tokens is not involved, and importantly, fuses and projects the outputs from independent branches to yield more expressive knowledge representation, as can be seen from our experimental comparisons.
Figure 2: Illustrative example of our four token-splitting strategies in action, showing how input tokens are partitioned and processed under each method.
3 Methodology
Figure 1 illustrates the overall architecture of our parallel hybrid network, FlowHN. It is composed of a cascade of parallel blocks, each of which contains two branches operating side by side: a SSM branch and an Attention branch. To allow the model to adaptively balance computation load between these two independent processing pathways, we introduce a dynamic token-splitting mechanism that distributes the input tokens between the parallel branches. While there is some information loss caused by the fact that components process the sequence not as a whole but in disjoint subsets, our split strategies are geared towards mitigating such loss. We explore three token splitting and allocation strategies. Each successor is an improvement over its predecessor in terms of load balancing between the two branches, as summarized in Figure 2. In the sections that follow, we describe in detail how each strategy partitions and routes tokens, how the branches process their respective input subsets, and how their outputs are subsequently fused and projected to improve the expressivity and yield the final representation from the parallel block.
3.1 No_Split Token Allocation Strategy
In this strategy called No_Split (Figure 2), the entire input token sequence is concurrently routed to both the SSM and Attention branches without any split. To reconcile the inherently divergent output formats of these pathways, we integrate a Token Fusion sub-module (depicted in blue in Figure 1) that serves as an information aggregator. Concretely, the dual-branch outputs are concatenated (not averaged as in Hymba (Dong et al. (2024))) and then passed through a linear projection, yielding a fused representation of the prescribed dimensionality. This fusion operation produces a semantically enriched embedding that retains complementary features from both branches while promoting feature interplay across modalities. Consequently, the network learns to optimally integrate these heterogeneous features, thereby enhancing the expressiveness and robustness of the resulting representation. As can be seen from our comparison results in Table  1, our fusion operation yields comparable (at times better) representation fidelity to that obtained through the use of meta-tokens in Hymba, and at much lower computational costs.
3.2 Alternating_Equal_Split Token Allocation Strategy
In this strategy called AE_Split, (Figure 2), the input token sequence is evenly partitioned between the SSM and Attention branches to halve the computational burden of each branch and accelerate the overall processing. Because each branch only sees half of the input in any given block, potentially relevant contextual information residing in the complementary half would be missed. To address this, we employ an alternating-half scheme whereby, in the subsequent block, each branch processes the opposite half of the sequence that it skipped previously. This “Alternative” scheduling thus ensures that over two consecutive block, both branches attend to all tokens, mitigating information loss while retaining per-block speed gains. Nevertheless, the faster branch must still idle until its slower counterpart completes its computation before the fused representation can be passed onward. This causes a synchronization bottleneck that limits the computational efficiency of the model.
Algorithm 1 FLOP-Aware Circulating Split (FAC_Split) strategy
1:Input: T = [$t_{1}$, $t_{2}$, …, $t_{L}$]
2:Output: Processed results from SSM and Attention branches
3:L: Number of tokens in the input sequence
4:T: Sequence of input tokens
5:F_a: Number of Attention FLOPs
6:F_s: Number of SSM FLOPs
7:block_index Number specifying block index
8:$block\_size\leftarrow\frac{L}{\frac{F_{s}}{F_{a}}+1}$triangleright SSM block processes $\frac{L}{\frac{F_{s}}{F_{a}}+1}$ tokens
9:$block\_index\leftarrow 0$
10:for $i=0$ to N-1 do $\triangleright$ N is Number of Parallel Blocks
11:     $start\_index\leftarrow(block\_index)\times block\_size$
12:     $end\_index\leftarrow(block\_index+1)\times block\_size$
13:     if $end\_index<L$ then
14:         $T_{SSM}\leftarrow T[start\_index:end\_index)$triangleright block_size tokens to SSM
15:         $T_{Attention}\leftarrow T[:start\_index\&end\_index:]$triangleright Rest of the tokens to Attention
16:         $SSM_{Output}\leftarrow SSM(T_{SSM})$
17:         $Attention_{Output}\leftarrow Attention(T_{Attention})$
18:         Merge $SSM_{Output}$ and $Attention_{Output}$triangleright Merge results of SSM and Attention
19:         $block\_index\leftarrow block\_index+1$
20:     else
21:         $block\_index\leftarrow 0$
22:     end if
23:end for
3.3 Flop_Aware_Split Token Allocation Strategy
This token allocation strategy called FA_Split introduces a FLOP-aware token-splitting mechanism to mitigate the synchronization bottleneck inherent in alternating branch design of the earlier strategy. Tokens are allocated to the SSM and Attention branches according to their respective FLOPs per token requirement. A smaller subset of the input sequence is dynamically assigned to the branch with higher FLOPs per token requirement, and the remaining larger subset to the other branch. This helps equitably divide the computational workload. FA_Split synchronizes branch completion times and thereby approximates maximally parallel execution, resulting in improved TPS and memory efficiency. Despite these advantages, this strategy, while preserving the original input order, inevitably processes disjoint subsets of tokens in each branch, thus not facilitating cross-branch information sharing, so essential for increasing representation expressivity. As shown in Figure 2, this hard split can lead to missing contextual cues from other branches, which can limit the model’s ability to capture global dependencies.
3.4 FLOP_Aware_Circulating_Split Token Allocation Strategy
With the goal of overcoming the limitations of previous modes, we propose a FLOP-Aware Circulating Split strategy (FAC_Split), that dynamically distributes tokens between the SSM and the Attention based on their relative computational costs. As illustrated in Figure 2, we first measure the FLOPs of each branch (SSM versus Attention) and then allocate tokens inversely proportional to those costs: the slower, higher-FLOP cost branch receives fewer tokens, while the faster, lower-FLOP cost branch receives more. For example in Figure 2, when the Attention’s FLOPs costs are twice those of the SSM, it processes only half that of the other branch in each parallel block. For example, in the first block, tokens 1–4 go to the SSM and tokens 5–6 to the Attention; in the next block, tokens 1–2 and 5–6 go to the SSM (with tokens 3–4 going to the Attention); and in the third block, tokens 1–2 go to the Attention and tokens 3–6 to the SSM. This cyclic reassignment ensures that, over the sequence of parallel blocks, both branches very likely see every token at least once. By balancing per-branch workloads, this token splitting mode achieves higher TPS and better memory efficiency than static splits, and by rotating token assignments, it promotes richer, more diverse feature learning without introducing synchronization bottlenecks. Algorithm 1 details the token-feeding procedure of FAC_Split strategy.
4 Experiments
4.1 Settings
Datasets. All models are trained on SlimPajama-6B, a 6 billion-token English subset sampled from the first chunk of the deduplicated SlimPajama-627B Soboleva et al. (2023) corpus that preserves its original domain mix (54.1% CommonCrawl, 28.7% C4, 4.2% GitHub code, 3.7% books, 3.4% arXiv, 3.1% Wikipedia, 2.8% StackExchange), using the publicly available Cerebras preprocessing scripts.
Baseline Models. To assess the effectiveness of FlowHN, we benchmarked it against several state-of-the-art baselines—including LLaMA, Mamba-1, Mamba-2, Jamba, Sequential Hybrid Model (SHM), and Hymba—across three model scales: 135M, 350M, and 1B parameters. SHM serves as the sequential counterpart to our proposed FlowHN, interleaving LayerNorm + SSM + LayerNorm + MLP with LayerNorm + Attention + LayerNorm + MLP in sequence.
4.2 Implementation details
All experiments were conducted on a single NVIDIA Tesla V100 PCIe GPU with 32GB of memory. All models were trained using a single GPU. The average training times for models with 135M, 350M, and 1B parameters were around 17 hours, 3 days, and 11 days, respectively. Due to limited resources, we were unable to train our proposed model, FlowHN, on massive datasets like those used for large-scale models such as LLaMA, Mamba, Jamba, and Hymba. For a fair and rigorous comparison among all state of the art models, we trained all models, FlowHN and the baselines, from scratch under identical experimental conditions, even if these conditions may not be individually optimal, and not for FlowHN either. Specifically, each configuration was trained using exactly 1,024 tokens. For the 135M and 350M parameter models, we used a batch size of 16 with gradient accumulation over 8 steps, an initial learning rate of $3\times 10^{-4}$, weight decay of 0.1, and AdamW with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. We applied a cosine learning-rate schedule with 10% linear warmup and trained for a total of 1B tokens. For the 1B parameter models, we reduced the batch size to $4$ and increased gradient accumulation to 16 to accommodate memory constraints, set the learning rate to $1\times 10^{-4}$, maintained the $0.1$ weight decay, and used AdamW with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ under the same cosine schedule and 10% warmup, training across 2B tokens.
4.3 Results and Discussion
Table 1 presents the performance of FlowHN with the four token allocation strategies and three model sizes (135M, 350M, and 1B parameters), compared against Llama, Mamba-1, Mamba-2, Jamba, the and the parallel hybrid model Hymba. This table also presents a comparison of the evaluated models in terms of Tokens per Second (TPS), Model FLOPs Utilization (MFU), and accuracy. MFU Chowdhery et al. (2023) quantifies the efficiency with which the available computational resources (e.g., GPUs or TPUs) are utilized during both the forward and backward passes to perform floating-point operations—an essential component of neural network computations. A higher MFU indicates better utilization of hardware capabilities, which can translate to faster training or inference and potentially improved model performance. MFU can be computed using either of the following equivalent equations:
$MFU$ $=\frac{\text{FLOPs/Iter}\times\text{Iter/Sec}}{\text{Device peak FLOPs/Sec}}$ $MFU$ $=\frac{\text{Token/Sec}\times\text{FLOPs/Token}}{\text{Device peak FLOPs/Sec}}$ (1)
Our evaluation results show that our proposed FlowHN framework achieves a much better balance between computational efficiency and downstream task accuracy at every model scale we tested. For the 135M-parameter models, No_Split attains an average accuracy of 38.85 %, outperforming the strongest baseline by nearly one full percentage point. It also sustains one of the highest TPS—26443, and also delivers an MFU of 21.99 %. At the 350M scale, No_Split again leads on accuracy (41.31 %), a 1.5 % improvement over Hymba, while nearly doubling on Hymba’s TPS performance, and also raising MFU to 25.12 %. Even at the one billion-parameter level, FlowHN variants continue to scale efficiently: FAC_Split reaches the maximum MFU of 47.63 %, compared to Hymba’s 33.16 %, demonstrating that our split based token distribution harnesses hardware capacity better as model size grows, with only a very slight lowering of accuracy.
Table 1: Comparative analysis of TPS, MFU, and accuracy across various state-of-the-art models, including our proposed FlowHN, evaluated at three model scales: 135M, 350M, and 1B parameters.
Model Name Tokens TPS MFU arc_e arc_c wino piqa hella boolq Avg
135M
Llama 1B 21353 22.33 27.23 22.95 47.97 52.42 25.12 41.24 36.15
Mamba-1 1B 11854 15.35 26.85 24.66 49.51 53.02 26.03 40.99 36.85
Mamba-2 1B 10094 13.29 28.16 23.89 48.54 53.13 25.45 41.08 36.71
Jamba 1B 18324 13.83 27.78 22.44 37.81 25.79 53.02 49.35 36.03
SHM 1B 25900 22.75 29.67 22.78 40.72 26.56 55.25 50.65 37.96
Hymba 1B 9393 12.21 29.88 23.89 49.03 55.04 25.56 44.57 37.99
No_splitours 1B 26443 21.99 29.46 24.23 49.92 56.02 27.00 43.44 38.85
AE_Splitours 1B 33846 22.80 26.22 23.81 48.62 52.59 25.95 40.81 36.33
FA_Splitours 1B 34348 23.02 25.29 23.98 47.81 50.57 25.26 38.48 35.23
FAC_Splitours 1B 34430 23.07 28.75 23.29 48.95 54.33 26.80 42.32 37.41
350M
Jamba 1B 13202 20.03 29.04 22.61 52.14 26.71 56.78 51.22 39.75
SHM 1B 8547 24.45 29.17 23.55 58.52 26.84 55.74 50.49 40.71
Hymba 1B 4452 14.89 30.58 24.49 49.43 54.22 26.12 53.82 39.78
No_splitours 1B 8587 25.12 30.83 25.88 49.27 57.59 27.03 57.23 41.31
AE_Splitours 1B 13697 25.18 27.44 23.46 49.51 55.42 25.71 44.37 37.65
FA_Splitours 1B 14246 26.41 25.42 22.53 48.54 53.08 26.01 40.81 36.07
FAC_Splitours 1B 14385 26.64 29.34 25.77 47.97 56.29 27.44 56.56 40.56
1B
Jamba 2B 9758 40.63 34.68 23.29 49.51 58.25 26.95 56.78 41.57
SHM 2B 4923 47.35 35.56 23.89 49.76 57.49 27.44 58.25 42.06
Hymba 2B 3283 33.16 37.04 21.97 49.27 59.39 27.18 60.11 42.41
No_splitours 2B 4325 45.02 36.15 22.95 49.51 59.39 28.27 61.46 42.95
AE_Splitours 2B 7752 46.16 33.38 22.53 49.19 55.42 26.80 53.72 40.17
FA_Splitours 2B 7901 47.48 32.03 23.46 49.43 54.16 25.82 46.10 38.50
FAC_Splitours 2B 7926 47.63 36.07 23.89 50.97 57.59 26.54 56.78 41.97
Figure 3 shows the comparison graph of MFU and TPS between our proposed FlowHN and three state-of-the-art hybrid baselines at three model scales: 135M, 350M, and 1B parameters. Across all sizes, the different token allocation strategies that resort to splitting achieve the highest TPS, with peaks of approximately 34430 TPS at 135M and sustaining over 7926 TPS at 1B. The sole exception is Jamba-1B, which achieves a somewhat higher TPS than FlowHN-1B. This advantage stems from Jamba-1B’s Mixture of Expert (MoE) design: by replacing every other feed-forward block with an MoE layer comprising eight experts, of which only two are active per token, the model maintains a total parameter count of 1B but effectively utilizes only 483M active parameters. The resulting 50% reduction in active parameters yields faster token processing, at the cost of somewhat lower MFU compared to our FlowHN-1B.
Figure 3 further reveals that FlowHN consistently achieves the highest MFU at all three scales (135M, 350M, and 1B parameters), with one exception: at the 350M scale, LLaMA slightly surpasses FlowHN. This discrepancy arises because LLaMA-350M incurs 27.46T FLOPs/Iter versus FlowHN’s 19.24T, and despite FlowHN’s higher TPS (14385 TPS vs. 9064 TPS), the TPS uplift is insufficient to fully offset the lower compute density—resulting in marginally lower MFU. In contrast, at 135M and 1B scales the FLOPs/Iter gap between LLaMA and FlowHN narrows, so FlowHN’s TPS gains more than compensate, yielding superior MFU across those settings.
Figure 3 also reflects the underlying scaling laws: as model size increases, the total number of floating-point operations per iteration (FLOPs/Iter) grows roughly in proportion to the number of parameters, which drives MFU upward. Conversely, deeper and wider layers incur greater data movement and inter-GPU communication overhead, thereby reducing TPS. According to Equation 1, since the denominator (theoretical peak FLOPs) remains constant while FLOPs per step increase more than the corresponding drop in Iter/Sec, MFU necessarily improves with scale.
Figure 3: Effect of model scale on TPS and MFU. Bars (left axis) show TPS, while lines (right axis) trace MFU for each model at 135M, 350M, and 1B parameters.
Our results in Figure 4 demonstrate the performance of our proposed parallel hybrid model variants relative to state-of-the-art hybrid architectures across three parameter scales. The No_split mode consistently delivers the largest performance improvements at all scales. FAC_Split remains competitive with existing hybrids, whereas FA_Split underperforms on all scales, its token-splitting procedure introduces discontinuities that hinder information flow and yields only modest gains at the 1B parameter scale. FAC_Split reduces the gap to full-parallel performance, indicating that cyclically reordering tokens so that both Attention and SSM sub-modules observe every token at least once provides benefits, although it still falls short of granting each sub-module complete context.
Figure 4: Accuracy as a function of model scale (135 M, 350 M, and 1 B parameters) comparing our approach against state-of-the-art hybrid models.
5 Conclusion
In this work, we introduce the FlowHN, a novel architecture that unifies attention and SSMs in a balanced, parallel fashion. By leveraging parallel processing and efficient FLOP aware token distribution, FlowHN ensures that both attention and SSM sub-modules contribute equitably to each block’s workload and addresses the TPS-latency trade-off inherent in traditional hybrid sequence models. Our extensive experiments across multiple model scales demonstrate that FlowHN consistently outperforms existing hybrid parallel architectures, achieves up to 4×4$\times$4 × relative gains in Token per Second and delivers up to 2×2$\times$2 × gain in MFU. The success of FlowHN underscores the potential of parallel hybrid integration in large-scale AI systems. Future work will explore the scalability of FlowHN to even larger models. We believe that FlowHN paves the way for more efficient and scalable architectures in AI technologies.
Limitations: The FlowHN has been primarily validated on autoregressive language modeling tasks, such as commonsense reasoning and question answering. Its performance across other NLP tasks, including machine translation, summarization, and recall-intensive applications, remains to be fully explored. Currently, FlowHN employs token-splitting based on computational workload; however, future work could investigate dynamic routing mechanisms that consider both computational efficiency and the relevance or "helpfulness" of each branch for processing specific tokens. Addressing these limitations will be helpful for enhancing the model’s robustness and generalization capabilities in our future work.
Broader Impact: This work proposes a FLOP-aware dynamic token splitting strategy that achieves significant speedups while not losing much on model accuracy. The resulting efficiency improvements have the potential to reduce per-token energy consumption, enable broader access to large-scale models on resource-constrained devices, and facilitate real-time AI applications. However, increased computational efficiency may also accelerate the automation of tasks, potentially requiring re-skilling in certain sectors.
Appendix A Technical Appendices and Supplementary Material
Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline. There is no page limit for the technical appendices.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See the Abstract.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Section 5
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: See Section 4.3
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See Section 4.2
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: Publicly available dataset has been used. Code will be released upon publication or provided upon request.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See Section 4.2
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: No further statistical experiments were deemed necessary, as the reported TPS and MFU results were obtained through multiple runs and consistently fell within the same range, demonstrating their stability and reliability.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Section 4.2
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, we confirm.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: See Section 5
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: There are no such problems in our task.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: See Section 4.1
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We use existing public benchmark datasets for experiments.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No such experiments or research are involved in our work
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No such experiments or research are involved in our work.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: It hasn’t been used as an important, original, or non-standard component of the core method.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 However, increased computational efficiency may also accelerate the automation of tasks, potentially requiring re-skilling in certain sectors.
-Appendix A Technical Appendices and Supplementary Material
-Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline.
-There is no page limit for the technical appendices.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: See the Abstract.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: See Section 5
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: See Section 4.3
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: See Section 4.2
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [No]
-  24. Justification: Publicly available dataset has been used. Code will be released upon publication or provided upon request.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: See Section 4.2
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [No]
-  34.
-Justification: No further statistical experiments were deemed necessary, as the reported TPS and MFU results were obtained through multiple runs and consistently fell within the same range, demonstrating their stability and reliability.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: In Section 4.2
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: Yes, we confirm.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: See Section 5
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: There are no such problems in our task.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: See Section 4.1
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: We use existing public benchmark datasets for experiments.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: No such experiments or research are involved in our work
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: No such experiments or research are involved in our work.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: It hasn’t been used as an important, original, or non-standard component of the core method.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.